OpenStudy (anonymous):

@Hero One more? :) A biologist catches 46 fish in a pond, tags them, and then releases them back into the pond. On a later visit to the pond, the biologist catches 20 fish, of which 8 are tagged. Which is the best estimate for the number of fish in the pond? A. 66 B. 115 C. 368 D. 920 IDK how to do this one. Explain? :)

OpenStudy (anonymous):

This tells me that the 46 she tagged represents 8/20 ths of the total amount of all fish or 2/5ths of the total 46 = 2/5 x x = 5/2 * 46 or about 2.5 times 46 I'd go with 115

OpenStudy (anonymous):

@Hero Is it 115?

hero (hero):

It is correct, but @BangkokGarrett should make sure you understand it.

hero (hero):

My explanation of it would be a bit different though.

OpenStudy (anonymous):

I don't. Can u explain @Hero ?

OpenStudy (anonymous):

It is a little confusing. She tags the fish. They then swim around and get all mixed up, right? So when she finds that 8 of a random sample of 20 are tagged...that tells her that 8/20 (that's 40%) of the fish are tagged So the question boils down to 46 is 40% of what number.

OpenStudy (anonymous):

Tag 46 fish and let em go. Do an experiment to figure out what % of the fish are tagged. The experiment tells us 8/20ths or 40%...are now tagged 46 represents 40% of all the fish. 46 is 40% of what? Use your favorite method to answer that.

OpenStudy (anonymous):

Feel free to ask questions about the parts you don't understand

hero (hero):

I have a different approach to this. I will post it soon. It is a more intuitive approach.

OpenStudy (anonymous):

I GOT A 100!! THANKS GUYS!! :) Though I would still like to know how you do it @Hero

OpenStudy (anonymous):

You're welcome